OpenStudy (joannablackwelder):

In a class of 80 students, 40 bring their textbooks to class, 20 bring some sort of beverage to class, 60 bring calculators to class, while 5 students bring nothing to class. If 30 students bring both their textbooks and calculators to class, how many students bring only a beverage to class? (Hint: Determine the number of students that bring a textbook or a calculator (or both) to class.)

OpenStudy (joannablackwelder):

@nincompoop Can you help?

OpenStudy (kropot72):

P(T) = 0.5 P(B) = 0.25 P(C) = 0.75 \[\large P(T \cap C)=0.375\] \[\large P(T \cup C)=P(T)+P(C)-P(T \cap C)\] \[\large =0.5+0.75-0.375=0.875=\frac{7}{8}\]

OpenStudy (joannablackwelder):

Wow, that is way easier than I was making it out to be. Thanks, @kropot72!

OpenStudy (haseeb96):

is the answer of 70 is correct

OpenStudy (kropot72):

So the number that bring either a textbook or a calculator is: \[\large 80\times\ \frac{7}{8}=70\] 5 students bring nothing to class. So how many students bring only a beverage to class?

OpenStudy (joannablackwelder):

5

OpenStudy (kropot72):

You are correct!

OpenStudy (joannablackwelder):

You are the best, thanks so much!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (haseeb96):

so the answer is 65 ?

OpenStudy (joannablackwelder):

No, I get the answer to be 5, and kropot72 confirmed.